# Automated Trustworthiness Oracle Generation for Machine Learning Text Classifiers

### v1

@@ 3. TOKI: Trustworthiness Oracle through Keyword Identification
As outlined in Section 2, a prediction is considered trustworthy if its reasoning is plausible. In text classification, a trustworthy prediction should rely on words semantically related to the predicted class. We argue that examining the distribution of words helps better recognise these semantically related words. Specifically, directly related words can act as anchors, and words indirectly related to the class are likely to form clusters around them. Following this intuition, we present Trustworthiness Oracle through Keyword Identification (TOKI). The key idea of TOKI is to use a list of keywords for each class to indicate what a text classifier should rely on to make predictions. To identify these keywords, TOKI selects clusters containing directly related words and their surrounding indirectly related words. A prediction is then deemed trustworthy if it mainly relies on the keywords of the predicted class. While the ultimate goal is to identify a complete list of keywords, this is impractical due to resource and computational constraints. Therefore, we identify only a partial list of keywords by extracting decision-contributing words from the classifier’s responses on the training data and applying clustering analysis to them. Since not all decision-contributing words are genuinely related to the class, clustering analysis also helps separate related words from unrelated ones. Figure 4 describes the process of TOKI, which consists of two pipelines: keyword identification and trustworthiness label computation.
[The process of TOKI including two pipelines]Fully described in the text.
@@ 3.1. Keyword Identification
Step 1: Explaining
To extract keywords, we focus on analysing the reasoning behind correct predictions, as incorrect ones might provide meaningless reasoning. Therefore, TOKI identifies the decision-contributing words from instances that have been predicted correctly in the training set. Let , , and represent the training set, the text classifier, and the true class, respectively. The explaining step is formalised in Equation 1, which returns a list of explanations for each class .
TOKI leverages the explanation method to measure the contribution of each word to a prediction. Several methods, such as LIME (Ribeiro et al., 2016), achieve this by locally approximating the model as an interpretable surrogate model. Other methods (Li et al., 2016) perturb input and evaluate model output changes. Another common way (Mohebbi et al., 2021) is to compute the gradient of the output with respect to the input. All these methods generate explanations for the prediction in the form of , a list of top decision-contributing words with their importance scores.
Step 2: Word pool construction
A word pool of decision-contributing words and their averaged importance scores across explanations is created for each class, as formalised in Equation 2.
@@
Step 3: Word clustering
The word pool of the class contains both keywords and unrelated words. To distinguish them and collect both directly related and indirectly related words, TOKI clusters the word pool , as formalised in Equation 3.
TOKI transforms the words in into embeddings using word embedding methods. Since the number of clusters is unknown, hierarchical clustering (Nielsen, 2016) is applied to group these word embeddings based on their cosine similarities. Clusters of the class are obtained by cutting the dendrogram, a hierarchical tree of relationships between the word embeddings, at the threshold distance dist.
Step 4: Keyword selection
The list of keywords is identified by selecting the clusters of words semantically related to the class name, as described in Equation 4.
@@
While dist needs manual configuration, relate can be automatically estimated by turning it into a binary classification problem. The key idea to determine relate is that related pairs of words can be found via synonyms. To accomplish this, the top 1,000 most common English words are taken from WordNet (Miller, 1995). Then, TOKI uses the Merriam-Webster dictionary to find all single-word synonyms of each word, resulting in approximately 32,000 pairs of related words. Another 32,000 random pairs of words are generated from WordNet to create a list of unrelated pairs. Finally, TOKI determines the value of relate through a binary search on the word embeddings of these lists. At each iteration, all 64,000 pairs are classified, with each pair of words considered as related if the cosine similarity between two word embeddings is higher than or equal to the current relate. The search stops when precision and recall for both related and unrelated classifications are balanced. The value of relate varies depending on the word embedding method, as different methods have their unique ways of embedding, thereby impacting the measurement of similarity between words.
@@ 3.2. Trustworthiness Label Computation
Word embeddings themselves can be biased due to their training data (Torregrossa et al., 2021), potentially affecting the ability to measure semantic relatedness. To mitigate this, TOKI applies ensemble learning (Leon et al., 2017) by employing different word embedding methods. We use both static embedding methods (Bengio et al., 2000; Pennington et al., 2014; Shazeer et al., 2016; Bojanowski et al., 2017), which produce a single output for each word and contextual embedding methods (Cer et al., 2018; Devlin et al., 2019), which generate different vectors for the same word based on its context. Each method has a different way of vectorizing words, resulting in different similarity measurements between them. In the first pipeline, this affects the computation of relate and the keyword identification. In the second pipeline, different embedding methods identify different similar words in the word pool, leading to different decisions about how the word is related to the class. Decisions made by all embedding methods are combined using plurality voting (Leon et al., 2017) in both pipelines.
@@ 4. Targeted Adversarial Attacks on Trustworthiness Vulnerabilities
We introduce a novel adversarial attack method guided by TOKI. The key idea is to weaken valid correlations by replacing words with similar ones that are weakly correlated to the original class, while strengthening spurious correlations by injecting unrelated words strongly correlated to other classes. Table 1 compares TOKI-guided attack method with existing adversarial attack methods, exemplified by the state-of-the-art A2T (Yoo and Qi, 2021). A2T uses the gradient of the loss to determine the substitution order of words based on their importance scores in the prediction. It then iteratively replaces each word with synonyms generated from a counter-fitted word embedding model (Mrkšić et al., 2016). A modification rate sets the maximum number of perturbations allowed in the original text. The generated texts are subsequently filtered to ensure part-of-speech consistency and semantic preservation by evaluating the cosine similarity between the sentence encodings of the original and perturbed texts. A2T has been validated and demonstrated to be a strong adversarial attack method (Zhou et al., 2024).
@@ 5. Trustworthiness Benchmark
Dong’s trustworthiness dataset
The human-based evaluation conducted by Dong (Dong, 2018) is one of the most relevant studies. This evaluation uses two datasets: a subset (Ribeiro et al., 2016) of the 20 newsgroups (20news) differentiating Christianity from Atheism, and movie reviews with sentiment labels. In the evaluation, text inputs, highlighting the top words identified by the explanation method, are shown to crowdworkers. They then guess the system’s output and state their confidence on a five-point Likert scale, ranging from “strongly disagree” to “strongly agree”. To derive a trustworthiness dataset, only answers, where the model correctly predicts the output, are selected from the original crowdworkers’ responses. The answers, where crowdworkers either guess incorrectly with high confidence (4–5) or correctly with low confidence (1–2), are deemed untrustworthy. Other answers, where the crowdworkers guess correctly with high confidence (4–5), are trustworthy. Trustworthiness labels from the crowdworkers for the same prediction are combined using plurality voting (Leon et al., 2017) to determine the final trustworthiness label.
@@
Dong’s (Dong, 2018) | movie | 311 | 47 | 358 | MLP | 0.832 | 10, 20 | importance |
20news | MLP | 0.939 | equivalent | |||||
Garg et al.’s (Garg et al., 2022) | CAMS | 1,206 | 739 | 1,945 | mentalbert-base-uncased | 0.397 | 10 | importance equivalent |
Mathew et al.’s (Mathew et al., 2021) | HateXplain | 3,002 | 304 | 3,306 | bert-base-uncased | 0.797 | 10 | importance equivalent |
@@
CAMS (Garg et al., 2022) is a corpus for classifying mental health issues from social media posts, while HateXplain (Mathew et al., 2021) is a dataset for hate speech detection. Both datasets provide ground-truth explanations for each instance. In CAMS, annotations highlight phrases used as inferences for predictions. In HateXplain, tokens are labeled as 0 or 1 to indicate whether they are part of the explanation. Trustworthiness datasets are created by explaining the models and comparing these explanations with the ground-truth ones. To measure the plausibility of the reasoning behind predictions, we follow Zini et al. (El Zini et al., 2022) and use explanation precision , where is the model explanation and is the ground-truth explanation. High precision suggests that the model explanation is unlikely to provide a word not in the ground-truth explanation. A prediction is then considered trustworthy if its explanation precision 0.5.
Importance-different trustworthiness dataset
@@
We create an additional trustworthiness dataset in a more straightforward manner, with the distribution of importance scores among words considered. Six datasets and corresponding models are selected from Huggingface: amazon_polarity, ag_news, rotten_tomatoes, yahoo_answers_topics, imdb, and emotion. We randomly sample 1,000 predictions and have three trained participants annotating their trustworthiness. For each prediction, the participants first guess the output of the text input. They then review the model prediction and its explanation generated by LIME (Ribeiro et al., 2016), a state-of-the-art method known for its effectiveness and faithfulness (Mariotti et al., 2024; Zhao et al., 2024). Finally, the participants label the prediction as trustworthy or untrustworthy. Only predictions where both the model and participants guess the correct output are considered. The trustworthiness label for each prediction is determined by plurality voting (Leon et al., 2017) based on all annotations. It is observed that the participants often deem a prediction untrustworthy if unrelated words receive significantly higher importance scores in its explanation. The final trustworthiness benchmark is summarised in Table 2.
Explanation method
@@
Three explanation methods are used to extract decision-contributing words.
Models under test
@@ 6.1. Baselines
Trustworthiness assessment
We adopt a naive approach, named Naive, which is based on model confidence. Naive considers a prediction untrustworthy if its confidence is lower than a threshold conf. For an ablation study, we use a TOKI’s variant, called TOKI (–K.I), which directly measures the relatedness between decision-contributing words and class names without identifying keywords.
Adversarial attack
@@ 6.3. Metrics
We frame trustworthiness assessment as a binary classification task with trustworthy and untrustworthy labels. The effectiveness is measured by: overall accuracy; precision, sensitivity, and F1-score, which evaluate performance in detecting trustworthy predictions; specificity measures performance in detecting untrustworthy predictions; and geometric mean (), which balances the classification performance on both labels. The efficiency of trustworthiness oracles is measured using their processing time in seconds.
@@ 6.4. Experimental Setup
We used six different word embedding methods for the word embedding ensemble model of TOKI.
NNLM (Bengio et al., 2000) learns embeddings and language models using a feedforward neural network.
GloVe (Pennington et al., 2014) generates word embeddings from corpus word-to-word co-occurrence matrices.
Swivel (Shazeer et al., 2016) generates low-dimensional embeddings from feature co-occurrence matrices.
FastText (Bojanowski et al., 2017) represents each word as a bag of character n-grams, capable of embedding misspelled, rare, or out-of-vocabulary words.
USE (Cer et al., 2018) encodes sentences into embeddings for transfer learning to other tasks.
Bert (Devlin et al., 2019) is the first deeply bidirectional, unsupervised language representation, pre-trained on a large text corpus to condition both left and right contexts of each word.
The trustworthiness benchmark described in Section 5 is used to evaluate trustworthiness oracles. For our datasets, 2,000 training instances are randomly sampled for explanations in the first pipeline of TOKI, while all training instances are used in the remaining datasets. Trustworthiness oracles employing LIME, omission, and gradient explanations are denoted by the suffixes “-lime”, “-omis”, and “-grad”, respectively. Regarding adversarial attacks, we implement our method using TextAttack (Morris et al., 2020), which already includes the implementation of A2T (Yoo and Qi, 2021), and use A2T’s default settings for both methods. We then attack random samples of up to 8,000 instances from each dataset listed in Table 2. All experiments are run 10 times on a Macbook M3 Pro with 12-core CPU, 18-core GPU, 18GB RAM, and 512GB SSD. The final results are averaged across these runs.

### v3

@@ 3. TOKI: Trustworthiness Oracle through Keyword Identification
As outlined in Section 2, a prediction is considered trustworthy if its reasoning is plausible. In text classification, a trustworthy prediction should rely on words semantically related to the predicted class. We argue that examining the distribution of words helps better recognise these semantically related words. Specifically, directly related words can act as anchors, and words indirectly related to the class are likely to form clusters around them. Following this intuition, we present Trustworthiness Oracle through Keyword Identification (TOKI). The key idea of TOKI is to use a list of keywords for each class to indicate what a text classifier should rely on for predictions. To identify these keywords, TOKI selects clusters containing directly related words and their surrounding indirectly related words. A prediction is then deemed trustworthy if it mainly relies on the keywords of the predicted class. While the ultimate goal is to identify a complete list of keywords, this is impractical due to resource and computational constraints. Hence, we identify only a partial list of keywords by extracting decision-contributing words from the classifier’s responses on the training data and applying clustering analysis to them. Since not all decision-contributing words are genuinely related to the class, clustering analysis also helps separate related words from unrelated ones. Figure 4 describes TOKI with two pipelines: keyword identification and trustworthiness label computation.
[The process of TOKI, including two pipelines]Fully described in the text.
@@ 3.1. Keyword Identification
Step 1: Explaining. To extract keywords, we focus on analysing the reasoning behind correct predictions, as incorrect ones might provide meaningless reasoning. Therefore, TOKI identifies the decision-contributing words from instances that have been predicted correctly in the training set. Let , , and represent the training set, the text classifier, and the true class, respectively. The explaining step is formalised in Equation 1, which returns a list of explanations for each class .
TOKI leverages the explanation method to measure the contribution of each word to a prediction. Several methods, such as LIME (Ribeiro et al., 2016), achieve this by locally approximating the model as an interpretable surrogate model. Other methods (Li et al., 2016) perturb input and evaluate model output changes. Another common way (Mohebbi et al., 2021) is to compute the gradient of the output with respect to the input. These methods explain the prediction in the form of , a list of top decision-contributing words with their importance scores.
Step 2: Word pool construction. A word pool of decision-contributing words and their averaged importance scores across explanations is created for each class, as formalised in Equation 2.
@@
Step 3: Word clustering. The word pool of the class contains both keywords and unrelated words. To distinguish them and collect both directly related and indirectly related words, TOKI clusters the word pool , as formalised in Equation 3.
Words in are transformed into embeddings by embedding methods. As the number of clusters is unknown, hierarchical clustering (Nielsen, 2016) is applied to group these word embeddings based on their cosine similarities. Clusters of the class are obtained by cutting the dendrogram, a hierarchical tree of relationships between the word embeddings, at the threshold distance dist.
Step 4: Keyword selection. The list of keywords is identified by selecting the clusters of words semantically related to the class name, as described in Equation 4.
@@
While dist needs manual configuration, relate can be automatically estimated by turning it into a binary classification problem. The key idea to determine relate is that related pairs of words can be found via synonyms. To accomplish this, the top 1,000 most common English words are taken from WordNet (Miller, 1995). Then, TOKI uses Merriam-Webster Dictionary (2002) to find all single-word synonyms of each word, resulting in approximately 32,000 pairs of related words. Another 32,000 random pairs of words are generated from WordNet to create a list of unrelated pairs. Finally, TOKI determines the value of relate through a binary search on the word embeddings of these lists. At each iteration, all 64,000 pairs are classified, with each pair of words considered as related if the cosine similarity between two word embeddings is higher than or equal to the current relate. The search stops when precision and recall for both related and unrelated classifications are balanced. The value of relate varies depending on the word embedding method, as different methods have their unique ways of embedding, thereby impacting the measurement of similarity between words.
@@ 3.2. Trustworthiness Label Computation
Word embeddings themselves can be biased due to their training data (Torregrossa et al., 2021), potentially affecting the ability to measure semantic relatedness. To mitigate this, TOKI applies ensemble learning (Leon et al., 2017) by employing different word embedding methods. We use both static embedding methods (Shazeer et al., 2016; Bojanowski et al., 2017), which produce a single output for each word and contextual embedding methods (Cer et al., 2018; Devlin et al., 2019), which generate different vectors for the same word based on its context. Each method has a different way of vectorizing words, resulting in different similarity measurements between them. In the first pipeline, this affects the computation of relate and the keyword identification. In the second pipeline, different embedding methods identify different similar words in the word pool, leading to different decisions about how the word is related to the class. Decisions made by all embedding methods are combined using plurality voting (Leon et al., 2017) in both pipelines.
Plurality voting is a simple yet effective voting method where each voter selects a single option, and the option with the most votes wins. In the first pipeline, a word receives a vote from an embedding method if the method identifies it as a keyword. The word is ultimately classified as a keyword if it receives the highest number of votes across all embedding methods. Similarly, in the second pipeline, a word is considered related to a class if it is identified as such by the highest number of embedding methods in the ensemble.
@@ 4. Targeted Adversarial Attacks on Trustworthiness Vulnerabilities
We introduce a novel adversarial attack method guided by TOKI. The key idea is to weaken valid correlations by replacing words with similar ones that are weakly correlated to the original class, while strengthening spurious correlations by injecting unrelated words strongly correlated to other classes. Table 1 compares TOKI-guided attack method with existing adversarial attack methods, exemplified by the SOTA A2T (Yoo and Qi, 2021).
A2T uses the gradient of the loss to determine the substitution order of words based on their importance scores in the prediction. It then iteratively replaces each word with synonyms generated from a counter-fitted word embedding model (Mrkšić et al., 2016). This embedding model is injected with antonymy and synonymy constraints into vector space representations to improve its ability to assess semantic similarity. For example, traditional word embedding models like GloVe (Pennington et al., 2014) consider “expensive” similar to its antonyms, “cheaper” and “inexpensive”. In contrast, the counter-fitted word embedding model prefers synonyms like “costly” and “overpriced”. A2T also sets a modification rate to constrain the maximum number of perturbations allowed. The generated texts are subsequently filtered to ensure part-of-speech consistency and semantic preservation by evaluating the cosine similarity between the sentence encodings of the original and perturbed texts. A2T has been validated and demonstrated as a strong adversarial attack method (Zhou et al., 2024).
@@ 5. Trustworthiness Benchmark
Dong’s trustworthiness dataset. The human-based evaluation conducted by Dong (2018) is one of the most relevant studies. This evaluation uses two datasets: a subset (Ribeiro et al., 2016) of the 20 newsgroups (20news) differentiating Christianity from Atheism, and movie reviews with sentiment labels. In the evaluation, text inputs, highlighting the top words identified by the explanation method, are shown to crowdworkers. They then guess the system’s output and state their confidence on a five-point Likert scale, ranging from “strongly disagree” to “strongly agree”. To derive a trustworthiness dataset, only answers, where the model correctly predicts the output, are selected from the original crowdworkers’ responses. The answers, where crowdworkers either guess incorrectly with high confidence (4–5) or correctly with low confidence (1–2), are deemed untrustworthy. Other answers, where the crowdworkers guess correctly with high confidence (4–5), are trustworthy. Trustworthiness labels from the crowdworkers for the same prediction are combined using plurality voting (Leon et al., 2017) to determine the final trustworthiness label.
@@
movie (Dong, 2018) | 311 | 47 | 358 | Multilayer perceptron (MLPs) | 0.832 | 10, 20 | importance | |
20news (Dong, 2018) | 0.939 | equivalent | ||||||
CAMS (Garg et al., 2022) | 1,206 | 739 | 1,945 | mentalbert-base- uncased | 0.397 | 10 | importance equivalent | |
HateXplain (Mathew et al., 2021) | 3,002 | 304 | 3,306 | bert-base-uncased | 0.797 | 10 | importance equivalent | |
Issues (Schulte et al., 2024) | 2,187 | 45 | 2,232 | sebert-base | 0.945 | 10 | importance different | |
@@
CAMS and HateXplain. CAMS (Garg et al., 2022) is a corpus for classifying mental health issues from social media posts, while HateXplain (Mathew et al., 2021) is a dataset for hate speech detection. Both datasets provide ground-truth explanations for each instance. In CAMS, annotations highlight phrases used as inferences for predictions. In HateXplain, tokens are labeled as 0 or 1 to indicate whether they are part of the explanation. Trustworthiness datasets are created by explaining the models and comparing these explanations with the ground-truth ones. To measure the plausibility of the reasoning behind predictions, we follow Zini et al. (2022) and use explanation precision , where is the model explanation and is the ground-truth explanation. High precision suggests that the model explanation is unlikely to provide a word not in the ground-truth explanation. A prediction is then considered trustworthy if its explanation precision 0.5.
Issues: A SE-specific dataset. We adopt the study of Schulte et al. (2024), which analyses explanations for the automated classification of bug and non-bug issues, a critical task in SE. These issues are reported in issue tracking systems, such as JIRA or Github. Each prediction is explained by explanation methods, such as LIME (Ribeiro et al., 2016). The authors then review the explanation based on multiple criteria, assigning a score of +1 if satisfied, 0 if neutral, and -1 if not. To assess a prediction’s trustworthiness, we focus on two criteria: related and unambiguous. Specifically, related means there is a clear relationship between the important words of the explanation and the prediction. On the other hand, unambiguous implies there are no words of mixed meaning or all words are used in their correct meaning with respect to the explanation. Predictions are deemed trustworthy if the average score of these two criteria is greater than 0. If the average score is lower than 0, they are deemed untrustworthy. We finally combine these decisions of all annotators using plurality voting (Leon et al., 2017) to determine the final trustworthiness label for each prediction.
@@
Importance-different trustworthiness dataset. We create an additional trustworthiness dataset that considers the distribution of importance scores of words. Six datasets and corresponding models are selected from Huggingface: amazon_polarity, ag_news, rotten_tomatoes, yahoo_answers_topics, imdb, and emotion. We randomly sample 1,000 predictions and have three trained participants annotating their trustworthiness. For each prediction, the annotators first guess the output of the text input. They then review the model prediction and its explanation generated by LIME (Ribeiro et al., 2016), a SOTA method known for its effectiveness and faithfulness (Mariotti et al., 2024; Zhao et al., 2024). Finally, the prediction is manually labelled as trustworthy or untrustworthy by annotators. Only predictions where both the model and annotators guess the correct output are considered. The trustworthiness label for each prediction is determined by plurality voting (Leon et al., 2017) based on all annotations. It is observed that the annotators often deem a prediction untrustworthy if unrelated words receive significantly higher importance scores in its explanation.
@@
Explanation method. Three methods are used to extract decision-contributing words.
Models under test. For our new datasets, popular fine-tuned models from Huggingface are chosen. For the remaining datasets, models provided by the authors are used. Table 2 summarises the final trustworthiness benchmark, the models under test, and their corresponding accuracies.
@@ 6.1. Baselines
We adopt a naive approach, named Naive, which assesses trustworthiness based on model confidence. Naive considers a prediction untrustworthy if its confidence is lower than a threshold conf. For ablation studies, we use a TOKI’s variant, called TOKI (–K.I), which directly measures the relatedness between decision-contributing words and class names without identifying keywords (Cho et al., 2024). We also compare TOKI-guided attack method with A2T (Yoo and Qi, 2021), a SOTA adversarial attack method that has been validated to be strong and effective (Zhou et al., 2024).
@@ 6.3. Metrics
We emphasise that our evaluations focus on the effectiveness of trustworthiness oracles in classifying predictions as trustworthy or not, rather than assessing the trustworthiness of ML models themselves. The effectiveness of trustworthiness oracles is determined by the alignment between their generated trustworthiness labels and human-annotated ones that reflect human perceptions of trustworthiness in the benchmark shown in Table 2. We frame trustworthiness oracles as binary classifiers with trustworthy and untrustworthy labels and evaluate their effectiveness using standard performance metrics for binary classification.
Accuracy: the proportion of predictions correctly labelled as trustworthy or untrustworthy.
Precision, sensitivity, and F1-score: the performance in detecting trustworthy predictions.
Specificity: the performance in detecting untrustworthy predictions.
Geometric mean (G-mean): the balance between the classification performance on trustworthy and untrustworthy predictions, computed as .
This usage of these metrics differs from that of model confidence in RQ1, which is assumed to be an unsuitable indicator of trustworthiness. G-mean serves as a balanced metric for evaluating classification performance on both trustworthy and untrustworthy predictions. A higher G-mean indicates better alignment between the trustworthiness benchmark and the oracles. Additionally, the efficiency of the trustworthiness oracles is assessed by their processing time in seconds.
@@ 6.4. Experimental Setup
We use six word embedding methods for TOKI’s word embedding ensemble model. ① NNLM (Bengio et al., 2000) learns embeddings and language models using a feedforward neural network. ② GloVe (Pennington et al., 2014) generates word embeddings from corpus word-to-word co-occurrence matrices. ③ Swivel (Shazeer et al., 2016) generates low-dimensional embeddings from feature co-occurrence matrices. ④ FastText (Bojanowski et al., 2017) represents each word as a bag of character n-grams, capable of embedding misspelled, rare, or out-of-vocabulary words. ⑤ USE (Cer et al., 2018) encodes sentences into embeddings for transfer learning to other tasks. ⑥ Bert (Devlin et al., 2019) is the first deeply bidirectional, unsupervised language representation, pre-trained on a large text corpus to condition both left and right contexts of each word.
We use the trustworthiness benchmark described in Section 5 to evaluate trustworthiness oracles. For our datasets, 2,000 training instances are randomly sampled for explanations in TOKI’s first pipeline, while all training instances are used in the remaining datasets. Trustworthiness oracles employing LIME, omission, and gradient explanations are denoted by the suffixes “-lime”, “-omis”, and “-grad”, respectively. Regarding adversarial attacks, we implement our method using TextAttack (Morris et al., 2020), which already includes A2T’s implementation (Yoo and Qi, 2021), and use A2T’s default settings for both. We then attack random samples of up to 8,000 instances from each dataset in Table 2. Experiments are run 10 times on a Macbook M3 Pro with 12-core CPU, 18-core GPU, 18GB RAM, and 512GB SSD. The final results are averaged across these runs.